# Revisiting the Acousto-Electric Effect

### v1

@@ II.3 Carrier density equation
$2.\partialderivative{n_{s}}{t}-\mu E_{0}\partialderivative{n_{s}}{x}=-\omega_{p}^{2}\tau_{c}\left(n_{s}+\frac{e}{\varepsilon}\partialderivative[2]{u}{x}\right)+D_{n}\partialderivative[2]{n_{s}}{x}$. (18)
@@ V.3 Thermo-acoustic effect and gain saturation
We can now see the gain saturation mechanism at work: Consider $v_{d}>v_{a}$ so the prefactor $(1-v_{d}/v_{a})$ in the loss expression is negative and there is gain. As a result of the gain $\alpha<0$ the acoustic intensity II will grow, and the correction to the drift velocity $v_{d}^{(1)}$ will become more negative, so the net drift velocity $v_{d}=v_{d}^{(0)}+v_{d}^{(1)}$ will become smaller and the gain will decrease according to the prefactor $(1-v_{d}/v_{a})$, that is, the gain will saturate as the acoustic intensity increases.
@@
${dI\over dx}=-\alpha I=-(\alpha_{0}I_{sat}){I/I_{sat}\over(1+I/I_{sat})}.$ (55)
@@ V.4 Current saturation
from which we see that for the case of gain $v_{d}^{(0)}>v_{a}$ the current decreases with increasing acoustic intensity II. The physical ideas used here to describe saturation are in keeping with the more general approach in the 1966 paper by Ozaki and Mikoshiba OzaMik66 .

### v2

@@ II.3 Carrier density equation
$2.\partialderivative{n_{s}}{t}-\mu E_{0}\partialderivative{n_{s}}{x}=-\omega_{p}^{2}\tau_{c}\left(n_{s}+\frac{e}{q}\partialderivative[2]{u}{x}\right)+D_{n}\partialderivative[2]{n_{s}}{x}$. (18)
@@ V.3 Thermo-acoustic effect and gain saturation
We can now see the gain saturation mechanism at work: Consider $v_{d}>v_{a}$ so the prefactor $(1-v_{d}/v_{a})$ in the loss expression is negative and there is gain. As a result of the gain $\alpha<0$ the acoustic intensity $I$ will grow, and the correction to the drift velocity $v_{d}^{(1)}$ will become more negative, so the net drift velocity $v_{d}=v_{d}^{(0)}+v_{d}^{(1)}$ will become smaller and the gain will decrease according to the prefactor $(1-v_{d}/v_{a})$, that is, the gain will saturate as the acoustic intensity increases.
@@
${dI\over dx}=-2\alpha I=-2(\alpha_{0}I_{sat}){I/I_{sat}\over(1+I/I_{sat})}.$ (55)
@@ V.4 Current saturation
from which we see that for the case of gain $v_{d}^{(0)}>v_{a}$ the current decreases with increasing acoustic intensity $I$. The physical ideas used here to describe saturation are in keeping with the more general approach in the 1966 paper by Ozaki and Mikoshiba OzaMik66 .